OpenStudy (coolshubhs):

48÷2(9+3) = ? explain your answer...

OpenStudy (anonymous):

288 Just Calculate It

OpenStudy (coolshubhs):

show how.... some say it's 2 and some say it's 288.. me too think it's 288 !

OpenStudy (anonymous):

Order of operations can help as well. 48/2(9+3)= parentheses first. 48/2(12)= then multiply 48/24= then divide =2 Parenthesis, Exponents, Multiply, Divide, Addition, Subtraction.

OpenStudy (coolshubhs):

@Johnelle .. in order of operations divide comes before multiply

OpenStudy (anonymous):

Division Goes Before Any Of It

OpenStudy (coolshubhs):

wtf is going on with it... 50% say it's 2 and 50% say 288 ... I'm gone mad... >_<

OpenStudy (anonymous):

PEMDAS?

OpenStudy (coolshubhs):

Use standard order of operations ..

OpenStudy (coolshubhs):

@Hero ... help !!!

OpenStudy (anonymous):

Its 288

OpenStudy (coolshubhs):

yes it is ... but :( many are saying it's 2 ...-.-

OpenStudy (coolshubhs):

just show the solution..

OpenStudy (anonymous):

All You Have To Do Is Calculate

OpenStudy (coolshubhs):

48÷2(9+3) 48÷2(12) now the next step?

OpenStudy (anonymous):

All You Had To Do Was 48÷2(9+3) =288 Do All Together

OpenStudy (anonymous):

48/2(12) 24(12) 288

OpenStudy (anonymous):

Yes